Observation of ν_{4} band transitions of NH_{4}^{+} and ND_{4}^{+}

\underline{Changhong Xia}, Maria M. Sanz and Stephen C. Foster, Department of Chemistry, Mississippi State University Mississippi State, MS 39762-9573.

The tetrahedral ammonium ion has two triply-degenerate infrared-active fundamentals. In 1983, Crofton and Oka; and Schafer,Begemann,Gudeman and Saykally observed the ν_{3} fundamental of NH_{4}^{+} in glow discharges. Several subsequent studies have been reported, but only one concerns the other IR-active fundamental. In this work, more than thirty new lines in the ν_{4} band of NH_{4}^{+} are reported, together with the first detection of the same band of ND_{4}^{+}. These ammonium ions were observed in a water-cooled hollow cathode discharge with a diode laser spectrometer. A program was written to fit the spectra to fourth order based upon the spherical top Hamiltonian of Robiette et al. and using an uncoupled basis set.